OpenStudy (anonymous):

PLEASEPLEASEPLEASE Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = -x2 + 2.5x + 1. What do the solutions of 0 = -x2 + 2.5x + 1 represent on the graph of the equation?

OpenStudy (anonymous):

Oh, I remember doing this... First plot \(Y_1=-x^2+2.5x+1\) and then \(Y_2=0\). Find the intersection of these points. (As an aside, on a TI-83/84, you do this using the Intersect command). As for what these solutions represents, it is: The solutions of the equation are the x-coordinates of the intersection points of \(Y_1\) and \(Y_2\).

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Is that the legit answer

OpenStudy (precal):

You could also, solve for the zeros. On the TI 84+ and 83+, you would do y1=-x2 + 2.5x + 1 2nd trace zero, mark the left and right bound and then it will give it to you.

OpenStudy (anonymous):

What makes you think I was making it up?